# Classification of same-gate quantum circuits and their space-time symmetries with application to the level-spacing distribution

### v1

@@ Abstract
Introduction.– Chaoticity and integrability are important theoretical notions. Integrability can allow for analytical results, while chaotic systems, in spite of unpredictability of trajectories, adhere to statistical laws. So-called toy models – the simplest models with a given property – play an important role. Classical single-particle $H$ in one dimension (1D) is always integrable; one needs at least a 3D phase space for chaos to be possible. This can also be achieved already in 1D Ott (1993) by a time-dependent $H(t)$, the simplest case being a “kicked” system of form $H(t)=\frac{p^{2}}{2}+V(q)\tau\delta_{\tau}(t)$, where $\delta_{\tau}(t)$ is a train of delta functions. A canonical example is the standard map Casati et al. (1979). Similar logic of taking a Floquet propagator (map) works for quantum systems as well. For instance, one can exemplify single-particle quantum chaos with a kicked top Haake (2010); Haake et al. (1987). Going to many-body quantum systems a plethora of possibilities opens up, one choice being a Floquet propagator that is a product of two simpler propagators, e.g. Ref. Prosen (1999). In light of experimental advances in noisy quantum simulations Bloch et al. (2008); Mi et al. (2021); Kim et al. (2023) it pays off to consider systems where the basic building block is a nearest-neighbor gate rather than the local Hamiltonian (applying two-site gates is also simpler in classical numerical simulations Schollwöck (2011)).
@@
We therefore focus on circuits where a one-unit-of-time Floquet propagator $F$ is composed of applying the same two-site unitary gate $V$ on all nearest-neighbor pairs of qubits in 1D – we call such systems simple circuits. Simple circuits have translational and temporal invariance, and, depending on the chosen gate $V$, span all dynamical regimes from integrability to chaos. Needles to say, such simple circuits have been extensively studied, a non-exhaustive list of only few of recent works includes Refs. Akila et al. (2016); Gritsev and Polkovnikov (2017); Ljubotina et al. (2019); Gopalakrishnan and Lamacraft (2019); Zhou and Nahum (2020); Klobas et al. (2021); Bertini et al. (2019); Chan et al. (2022); Nahum et al. (2022); Claeys et al. (2022); Žnidarič (2023). We show that any such circuit has a very simple form: it is a product of a local 2-site transformation and a free propagation (translation), and that there are only $N-1$ spectrally inequivalent classes.
@@
There are many indicators of quantum chaos with perhaps the most frequently used one being the so-called level spacing distribution (LSD) $P(s)$ of nearest-neighbor eigenenergy spacing $s$ Haake (2010). According to the quantum chaos conjecture Bohigas et al. (1984) Hamiltonian systems with a chaotic classical limit  [^1] For many-body quantum systems a word of caution is in place. Because the level spacing is exponentially small in $N$ the LSD probes “chaoticity” on an exponentially large timescale. There can be situations where the relevant physics is not chaotic at all despite the RMT LSD, see e.g. Ref. Santos (2004); Brenes et al. (2018) are expected to display $P(s)$ given by the random matrix theory (RMT) Mehta (2004). For Floquet systems checking for quantum chaos via $P(s)$ is even simpler: writing eigenvalues of $F$ as ${\rm e}^{{\rm i}\phi_{j}}$ the density of eigenphases $\phi_{j}$ should be uniform and therefore taking for $s=(\phi_{j+1}-\phi_{j}){\cal N}/2\pi$ there is no need for the unfolding that is required when studying spectra of $H$ Haake (2010). It is therefore rather surprising that, while there are hundreds of papers using $P(s)$ to study chaoticity in many-body Hamiltonian systems, there are essentially none studying LSD in simple (same-gate) quantum circuits (exceptions are recent Refs. Duh ; Huang et al. (2023)). The reason is that, surprisingly, LSD for chaotic simple circuits seemingly does not adhere to the RMT expectation. In our paper we will show that the reason behind it is a space-time symmetry that all such circuits posses.
@@
Let us demostrate that by a simple generalized brick-wall (BW) circuit with 3 layers (Fig. 1 inset), where we translate each 2-site gate by 3 sites (instead of 2 as in BW). For periodic boundary conditions and $N$ divisible by $3$ the Floquet propagator can be written as $F=f_{2}f_{3}f_{1}$, where $f_{j}=\prod_{k=0}^{N/3-1}V_{j+3k,j+3k+1}$ is one layer beginning at site $j$ and $V_{i,j}$ denotes the unitary 2-site gate $V$ acting on qubits $i$ and $j$. Indices are taken modulo $N$, with sites $j=1,\dots,N$.
@@
We can also easily see (Fig. 1) that translating $F$ by 2 sites is the same as a shift in time by 1 layer. Denoting propagator from time $t_{1}$ to $t_{2}$ by $F(t_{1},t_{2})$, e.g., $F=F(0,1)$, the 3-layer BW circuits has a space-time symmetry $S^{-2}F(0,1)S^{2}=F(1/3,4/3)$ (application of each gate advances time by $1/N$). This symmetry is also reflected in the structure of $F$, which can be written as $F=S^{-4}f_{1}S^{4}S^{-2}f_{1}S^{2}f_{1}=S^{-6}(S^{2}f_{1})^{3}=S^{-6}\tilde{F}^{3}$. We now see where the crux of the problem lies. Since $F$ as well as $f_{1}$ has translational symmetry by 3 sites the momentum $k$ labeling eigenvalues of $S^{3}$ is a good quantum number. In each momentum eigenspace $S^{-6}$ is just an overall phase, and therefore $F$ is, up-to this irrelevant phase, equal to the 3rd power of $\tilde{F}=S^{2}f_{1}$. The quantum chaos conjecture should therefore be applied to $\tilde{F}$ rather than to $F$. Doing that one recovers perfect agreement with the RMT (Fig. 1). It also tells us that, provided $\tilde{F}$ and therefore the circuit is “chaotic”, the LSD of $F$ will be equal to that of the 3rd power of a CUE matrix which is equal to a direct sum of 3 independent CUE matrices, Eq. (9).
@@
Classification of simple circuits.– Having an $N$-site 1D system with periodic boundary conditions there are $N$ nearest-neighbor gates $V_{j,j+1}$ that can be ordered in $N!$ different ways (configurations) to make a one-step propagator $F$  [^2] Single-site Hilbert space can have any dimension (qudits) although we will frequently speak about qubits because all our numerical examples will be of that kind.. However, it is clear that many of those configurations have equal $F$ since 2-site gates acting on non-overlapping nearest neighbor sites commute. Furthermore, a lot of configurations lead to the same spectra, for instance, under cyclic permutations of gates spetra do not change Bensa and Žnidarič (2021). Because we want to study spectra of $F$ we will call two circuits equivalent if they have the same spectrum. It is clear that there are much less than $N!$ non-equivalent simple circuit classes. For open boundary conditions there is in fact just one class Bensa and Žnidarič (2021).
@@
Figure 2: Any same-gate nearest-neighbor circuit is spectrally equivalent to one of the canonical circuits $F_{q,r}$ shown here for $N=6$, see also Table 1.
For periodic boundary conditions this is not the case. One can show (see Appendix, Theorem S1) that there are $(N-1)$ different equivalence classes. A canonical representative circuit of a given equivalence class follows a similar logic as the 3-layer BW example in introduction: a canonical circuit is characterized by two integers $q$ and $r$, where the first layer of gates $f_{1}$ is made by repeatedly translating $V_{1,2}$ by $q$ sites ($q=3$ in the example), whereas $r$ determines the shift of the 2nd layer with respect to the 1st one ($r=2$ in the example). The following theorem embodies the precise statement.
@@
The proof can be found in Appendix C. While it is not constructive in the sense of providing an explicit procedure of transforming a given $F$ to its canonical form $F_{q,r}$, the transformation is in practice easily achieved by hand for small $N$, or one can in linear time calculate the invariant $p$ introduced in Lemma S1 in Appendix C, thereby obtaining the correct $(q,r)$. The integer invariant $p$ is defined for an alternative simple circuit representative of a given class, and characterizes the circuit as a concatenation of two staircase sections with opposite chirality, the length of the 2nd being $p$.
@@
The $(N-1)$ classes described in Theorem 1 account for all possible circuits of which the standard brick-wall with $(2,1)$, and the staircase Nahum et al. (2018); Lin et al. (2021); Bensa and Žnidarič (2021) (also called convolutional codes Poulin et al. (2009); Anikeeva et al. (2021)) with $(N,1)$ are just two cases. The allowed values of $(q,r)$ for few small $N$ are listed in Table 1, while their pictures are shown for $N=6$ in Fig. 2, and for $N=10$ in Fig. S1 in Appendix. Note that while the allowed set of $(q,r)$ for a given $N$ depends on the factors of $N$, the allowed $N$s for a given $(q,r)$ are simpler: taking any coprime $q>r$ a circuit is possible for all $N$ that are multiples of $q$.
@@
The set of allowed $(q,r)$ naturally splits into two categories. Because $q$ divides $N$, with the maximal value being $N$, one group is composed of the largest possible $q=N$, while the other has smaller $22\leq q\leq N/$ 2. Group (i) are generalized S circuits with $q=N$. Because the translation by $N$ is equivalent to the identity this group could be equivalently described by $(N,r)\equiv(r,0)$, that is, by a single integer $r$ that gives the shift of the next gate. As $r$ is coprime with $N$ all n.n. gates are obtained by a simple (ergodic) translation modulo $N$. Group (ii) can be viewed as generalized BW circuits and needs two integers. Because $q$ divides $N$, translation by $q$ alone is not ergodic and one needs subsequent layers characterized by $r$. Altogether one has a $q$-layer BW circuit, each layer consisting of $N/q$ gates.
@@
Figure 3: Level-spacing distribution $P(s)$ of $F_{q,r}$ (Eq. 2) in blue are not chaotic and are equal to a direct sum of $q$ CUE matrices (red curve, Appendix B), while the LSD of the root $\tilde{F}_{q,r}$ (Eq. 5, violet) agrees with the CUE RMT prediction (green curve). For $q\neq N$ we show data from the momentum eigenspace of $S^{q}$ with $k=0$, and averaging is done over $10-50$ circuits.
@@
For a generic gate $V$ the spectra of all $(N-1)$ propagators $F_{q,r}$ are different. If $V$ would be symmetric with respect to the exchange of the two qubits the circuits with $(q,r)$ and $(q,q-r)$ would have the same spectra (spatial reflection symmetry), and therefore one would have only $\lfloor\frac{N}{2}\rfloor$ different spectral classes Bensa and Žnidarič (2021). Theorem 1 also shows that for prime $N$ only generalized S circuits exist. For odd $N$ there are no standard BW circuits having $(2,1)$, but there are generalized BW circuits with $q>2$ (see Table 1). Note also that while for smaller $N<10$ possible circuits are straightforward generalizations of the S or BW configurations with left or right chirality, for larger $N$ less intuitive circuits are also possible. For instance, for $N=10$ one can have $(5,2)$ (see Fig. S1 in Appendix) that can be further compressed in time direction (e.g., all gates in the first two layers $f_{1}$ and $f_{3}$ commute), reducing the number of non-commuting layers from $q=5$ to just $3$. Each of those compressed layers has $2$ idle qubits (that are not acted upon), such that the compressed circuit $F_{5,2}^{t}$ has two separate lines of idle qubits, each of width $1$, slanted at $-45^{\circ}$. Integers $(q,r)$ therefore also determine the filling fraction, i.e., the number and the pattern of idle qubits in maximally compressed $F_{q,r}^{t}$ (see Appendix A). The only circuit with no idle qubits is the standard BW with $(2,1)$.
@@
Space-time symmetries.– In order to understand space-time symmetries of any simple circuit it suffices to study the canonical equivalence class representatives $F_{q,r}$. Denoting the inner term in Eq. (2) by $\tilde{F}_{q,r}$, calling it a root of $F_{q,r}$,
@@
This can be easily seen by rewriting the LHS of Eq. (8) as $S^{-qr}\left(S^{r}S^{-r}(S^{q}V_{1,2})^{N/q}S^{r}\right)^{q}$, which then equals to $S^{-qr}\left(S^{r}(S^{q}V_{1+r,2+r})^{N/q}\right)^{q}$. The final expression can be understood as a circuit beginning with the second layer, thus justifying the equality to the RHS of Eq. (8). Eq. (8) appeared in Ref. Huang et al. (2023) for the special case of $r=1$.
Level spacing statistics.– An immediate application of the above results is in quantum chaos for the statistics of spacings of closest eigenphases of $F$. Looking at the root connection in Eq. (6) and the fact that $S^{q}$ commutes with all terms, one can focus on a given common momentum eigenspace of $S^{q}$ with eigenvalues $e^{2\pi{\rm i}kq/N}$, $k\in\{0,1,\dots,N/q-1\}$. There $S^{-qr}$ is just an overall phase factor $e^{-2\pi{\rm i}qkr/N}$. Therefore $F_{q,r}$ is up-to this phase equal to an appropriate power of $\tilde{F}_{q,r}$. Eigenphases of $F_{q,r}$ are simple multiples of eigenphases of $\tilde{F}_{q,r}$ modulo $2\pi$, and therefore when trying to test the validity of the quantum chaos conjecture one must consider the level-spacing distribution of the root $\tilde{F}_{q,r}$ instead of the whole Floquet propagator $F_{q,r}$.
If the circuit is quantum chaotic one would therefore expect that the spectral statistics of the root $\tilde{F}_{q,r}$ (5) will adhere to the RMT theory. In particular, if the 2-site gate $V$ does not have any anti-unitary (time-reversal) symmetry, which is the case for our numerics where $V$ is randomly picked according to the Haar measure, the appropriate ensemble is the CUE. We can see in Fig. 3 that the LSD of $\tilde{F}_{q,r}$ indeed agrees with CUE Wigner surmise $P(s)=\frac{32}{\pi^{2}}s^{2}e^{-s^{2}4/\pi}$ for all canonical classes. Furthermore, the theorem by Rains Rains (2003) tells us what is the distribution of a power of a matrix from the CUE. For $M\in\mathrm{CUE}(\mathcal{N})$, where $\mathrm{CUE}(\mathcal{N})$ denotes a distribution of $\mathcal{N}\times\mathcal{N}$ CUE matrices, one has
$M^{q}\in\bigoplus_{j=1}^{q}\mathrm{CUE}\left(\left\lceil\frac{\mathcal{N}-j}{q}\right\rceil\right),$ (9)

### v2

@@ I Introduction
Chaoticity and integrability are important theoretical notions. Integrability can allow for analytical results, while chaotic systems, in spite of unpredictability of trajectories, adhere to statistical laws. So-called toy models – the simplest models with a given property – play an important role. Classical single-particle $H$ in one dimension (1D) is always integrable; one needs at least a 3D phase space for chaos to be possible. This can also be achieved already in 1D Ott (1993) by a time-dependent $H(t)$, the simplest case being a “kicked” system of form $H(t)=\frac{p^{2}}{2}+V(q)\tau\delta_{\tau}(t)$, where $\delta_{\tau}(t)$ is a train of delta functions. A canonical example is the standard map Casati et al. (1979). Similar logic of taking a Floquet propagator (map) works for quantum systems as well. For instance, one can exemplify single-particle quantum chaos with a kicked top Haake (2010); Haake et al. (1987). Going to many-body quantum systems a plethora of possibilities opens up, one choice being a Floquet propagator that is a product of two simpler propagators, e.g. Ref. Prosen (1999). In light of experimental advances in noisy quantum simulations Bloch et al. (2008); Mi et al. (2021); Kim et al. (2023) it pays off to consider systems where the basic building block is a nearest-neighbor gate rather than the local Hamiltonian (applying two-site gates is also simpler in classical numerical simulations Schollwöck (2011)).
@@
We therefore focus on circuits where a one-unit-of-time Floquet propagator $F$ is composed of applying the same two-site unitary gate $V$ on all nearest-neighbor pairs of qubits in 1D – we call such systems simple circuits. Simple circuits have translational and temporal invariance, and, depending on the chosen gate $V$, span all dynamical regimes from integrability to chaos. Needles to say, such simple circuits have been extensively studied, a non-exhaustive list of only few of recent papers includes Refs. Akila et al. (2016); Gritsev and Polkovnikov (2017); Ljubotina et al. (2019); Gopalakrishnan and Lamacraft (2019); Zhou and Nahum (2020); Klobas et al. (2021); Bertini et al. (2019); Chan et al. (2022); Nahum et al. (2022); Claeys et al. (2022); Žnidarič (2023). We show that any such circuit has a very simple form: it is a product of a local 2-site transformation and a free propagation (translation), and that there are only $N-1$ spectrally inequivalent classes.
@@
There are many indicators of quantum chaos with perhaps the most frequently used one being the so-called level spacing distribution (LSD) $P(s)$ of nearest-neighbor eigenenergy spacing $s$ Haake (2010). According to the quantum chaos conjecture Bohigas et al. (1984) Hamiltonian systems with a chaotic classical limit are expected to display $P(s)$ given by the random matrix theory (RMT) Mehta (2004). RMT LSD has also been observed in non-integrable generic systems without a classical limit, of which simple circuits are an example, where it is sometimes even taken as a defining property of quantum chaos  [^1] For many-body quantum systems a word of caution is in place. Because the level spacing is exponentially small in system size $N$ the LSD probes properties on an exponentially large timescale. There can be situations where the relevant physics is not chaotic at all despite the RMT LSD, see e.g. Refs. Santos (2004); Brenes et al. (2018). For Floquet systems checking for quantum chaos via $P(s)$ is even simpler: writing eigenvalues of $F$ as ${\rm e}^{{\rm i}\phi_{j}}$ the density of eigenphases $\phi_{j}$ should be uniform and therefore taking for $s=(\phi_{j+1}-\phi_{j}){\cal N}/2\pi$, where $\mathcal{N}$ is the Hilbert space dimension, there is no need for the unfolding that is required when studying spectra of $H$ Haake (2010). It is therefore rather surprising that, while there are hundreds of papers using $P(s)$ to study chaoticity in many-body Hamiltonian systems, there are essentially none studying LSD in simple (same-gate) quantum circuits (exceptions are recent Refs. Duh ; Huang et al. (2023)). The reason is that, surprisingly, LSD for chaotic simple circuits seemingly does not adhere to the RMT expectation. In our paper we will show that the reason behind it is a space-time symmetry that all such circuits posses.
@@
Let us demonstrate that by a simple generalized brick-wall (BW) circuit with 3 layers (Fig. 1 inset), where we translate each 2-site gate by 3 sites (instead of 2 as in BW). For periodic boundary conditions and $N$ divisible by $3$ the Floquet propagator can be written as $F=f_{2}f_{3}f_{1}$, where $f_{j}=\prod_{k=0}^{N/3-1}V_{j+3k,j+3k+1}$ is one layer beginning at site $j$ and $V_{i,j}$ denotes the unitary 2-site gate $V$ acting on qubits $i$ and $j$. Indices are taken modulo $N$, with sites $j=1,\dots,N$.
@@
We can also easily see (Fig. 1) that translating $F$ by 2 sites is the same as a shift in time by 1 layer. Denoting propagator from time $t_{1}$ to $t_{2}$ by $F(t_{1},t_{2})$, e.g., $F=F(0,1)$, the 3-layer BW circuits has a space-time symmetry $S^{-2}F(0,1)S^{2}=F(1/3,4/3)$ (application of each gate advances time by $1/N$). This symmetry is also reflected in the structure of $F$, which can be written as $F=S^{-4}f_{1}S^{4}S^{-2}f_{1}S^{2}f_{1}=S^{-6}(S^{2}f_{1})^{3}=S^{-6}\tilde{F}^{3}$. We now see where the crux of the problem lies. Since $F$ as well as $f_{1}$ have translational symmetry by 3 sites the momentum $k$ labeling eigenvalues of $S^{3}$ (also referred to as the quasi-momentum, since it takes a discrete set of values) is a good quantum number. In each momentum eigenspace $S^{-6}$ is just an overall phase, and therefore $F$ is, up-to this irrelevant phase, equal to the 3rd power of $\tilde{F}=S^{2}f_{1}$. The quantum chaos conjecture should therefore be applied to $\tilde{F}$ rather than to $F$. Doing that one recovers perfect agreement with the RMT (Fig. 1). It also tells us that, provided $\tilde{F}$ and therefore the circuit is “chaotic”, the LSD of $F$ will be equal to that of the 3rd power of a CUE matrix which is equal to a direct sum of 3 independent CUE matrices, Eq. (9).
@@
II Classification of simple circuits
Having an $N$-site 1D system with periodic boundary conditions there are $N$ nearest-neighbor gates $V_{j,j+1}$ that can be ordered in $N!$ different ways (configurations) to make a one-step propagator $F$  [^2] Single-site Hilbert space can have any dimension (qudits) although we will frequently speak about qubits because all our numerical examples will be of that kind.. However, it is clear that many of those configurations have equal $F$ since 2-site gates acting on non-overlapping nearest neighbor sites commute. Furthermore, a lot of configurations lead to the same spectra, for instance, under cyclic permutations of gates spetra do not change Bensa and Žnidarič (2021). Because we want to study spectra of $F$ we will call two circuits equivalent if they have the same spectrum. It is clear that there are much less than $N!$ non-equivalent simple circuit classes. For open boundary conditions there is in fact just one class Bensa and Žnidarič (2021).
@@ II Classification of simple circuits
Figure 2: Any same-gate nearest-neighbor circuit is spectrally equivalent to one of the canonical circuits $F_{q,r}$ shown here for $N=6$, see also Table 1. Dashed rectangles denote $f_{1}$, one layer occurring in the root $\tilde{F}_{q,r}$ in Eq. (5). Circuits (a), (b), and (c) are generalized BW while (d) and (e) are generalized S. Circuits (b) and (c), as well as (d) and (e), are chiral pairs.
For periodic boundary conditions this is not the case. One can show (see Appendix, Theorem 2) that there are $(N-1)$ different equivalence classes. A canonical representative circuit of a given equivalence class follows a similar logic as the 3-layer BW example in the introduction: a canonical circuit is characterized by two integers $q$ and $r$, where the first layer of gates $f_{1}$ is made by repeatedly translating $V_{1,2}$ by $q$ sites ($q=3$ in the example), whereas $r$ determines the shift of the 2nd layer with respect to the 1st one ($r=2$ in the example). The following theorem embodies the precise statement.
@@
The proof can be found in Appendix C. While it is not constructive in the sense of providing an explicit procedure of transforming a given $F$ to its canonical form $F_{q,r}$, the transformation is in practice easily achieved by hand for small $N$, or one can in linear time calculate the invariant $p$ introduced in Lemma 1 in Appendix C, thereby obtaining the correct $(q,r)$. The integer invariant $p$ is defined for an alternative simple circuit representative of a given class, and characterizes the circuit as a concatenation of two staircase sections with opposite chirality (Fig. 5), the length of the 2nd being $p$.
@@
The $(N-1)$ classes described in Theorem 1 account for all possible circuits of which the standard brick-wall with $(2,1)$, and the staircase Nahum et al. (2018); Lin et al. (2021); Bensa and Žnidarič (2021) (also called convolutional codes Poulin et al. (2009); Anikeeva et al. (2021)) with $(N,1)$ are just two cases. The allowed values of $(q,r)$ for few small $N$ are listed in Table 1, while their pictures are shown for $N=6$ in Fig. 2, and for $N=10$ in Fig. 4 in Appendix. Note that while the allowed set of $(q,r)$ for a given $N$ depends on the factors of $N$, the allowed $N$s for a given $(q,r)$ are simpler: taking any coprime $q>r$ a circuit is possible for all $N$ that are multiples of $q$.
@@
The set of allowed $(q,r)$ naturally splits into two categories. Because $q$ divides $N$, with the maximal value being $N$, one group is composed of the largest possible $q=N$, while the other has smaller $22\leq q\leq N/$ 2. Group (i) are generalized S circuits with $q=N$. Because the translation by $N$ is equivalent to the identity this group could be equivalently described by $(N,r)\equiv(r,0)$, that is, by a single integer $r$ that gives the shift of the next gate. As $r$ is coprime with $N$ all n.n. gates are obtained by just this translation modulo $N$. Group (ii) can be viewed as generalized BW circuits and needs two integers. Because $q$ divides $N$, translation by $q$ alone does not generate all n.n gates and one needs subsequent layers characterized by $r$. Altogether one has a $q$-layer BW circuit, each layer consisting of $N/q$ gates.
@@
Figure 3: Level-spacing distribution $P(s)$ of eigenphases of $F_{q,r}$ (Eq. 2) in blue are not chaotic and are equal to a direct sum of $q$ CUE matrices (red curve, Appendix B), while the LSD of the eigenphases of root $\tilde{F}_{q,r}$ (Eq. 5, violet) agrees with the CUE RMT prediction (green curve). For $q\neq N$ we show data from the momentum eigenspace of $S^{q}$ with $k=0$, and averaging is done over $10-50$ circuits each having an independent 2-site Haar random gate $V$.
@@
For a generic gate $V$ the spectra of all $(N-1)$ propagators $F_{q,r}$ are different. If $V$ would be symmetric with respect to the exchange of the two qubits the circuits with $(q,r)$ and $(q,q-r)$ would have the same spectra (spatial reflection symmetry), and therefore one would have only $\lfloor\frac{N}{2}\rfloor$ different spectral classes Bensa and Žnidarič (2021). Theorem 1 also shows that for prime $N$ only generalized S circuits exist. For odd $N$ there are no standard BW circuits having $(2,1)$, but there are generalized BW circuits with $q>2$ (see Table 1). It is interesting to note that circuits with more complex multi-site update rules have been used before, for instance the $(3,1)$ case Pozsgay (2021) as well as $(4,1)$ has been used to construct integrable models Gombor and Pozsgay (2021) (although with a 3-site transformation). One interesting question is possible integrability of different canonical configurations for specific $V$. While for smaller $N<10$ possible circuits are straightforward generalizations of the S or BW configurations with left or right chirality, for larger $N$ less intuitive circuits are also possible. For instance, for $N=10$ one can have $(5,2)$ (see Fig. 4 in Appendix) that can be further compressed in time direction (e.g., all gates in the first two layers $f_{1}$ and $f_{3}$ commute), reducing the number of non-commuting layers from $q=5$ to just $3$. Each of those compressed layers has $2$ idle qubits (that are not acted upon), such that the compressed circuit $F_{5,2}^{t}$ has two separate diagonally slanted lines of idle qubits, each of width $1$. Integers $(q,r)$ therefore also determine the filling fraction, i.e., the number and the pattern of idle qubits in maximally compressed $F_{q,r}^{t}$ (see Appendix A). The only circuit with no idle qubits is the standard BW with $(2,1)$.
III Space-time symmetries
@@
In order to understand space-time symmetries of any simple circuit it suffices to study the canonical equivalence class representatives $F_{q,r}$. Denoting the inner term in Eq. (2) by $\tilde{F}_{q,r}$, calling it a root of $F_{q,r}$,
@@
This can be easily seen by rewriting the LHS of Eq. (8) as $S^{-qr}\left(S^{r}S^{-r}(S^{q}V_{1,2})^{N/q}S^{r}\right)^{q}$, which then equals to $S^{-qr}\left(S^{r}(S^{q}V_{1+r,2+r})^{N/q}\right)^{q}$. The final expression can be understood as a circuit beginning with the second layer, thus justifying the equality to the RHS of Eq. (8). Eq. (8) appeared in Ref. Huang et al. (2023) for the special case of $r=1$, along with a figure showing numerically computed LSD of $F_{q,r}$ for $q=2,r=1$.
IV Level spacing statistics
An immediate application of the above results is in quantum chaos for the statistics of spacings of closest eigenphases of $F$. Looking at the root connection in Eq. (6) and the fact that $S^{q}$ commutes with all terms, one can focus on a given common momentum eigenspace of $S^{q}$ with eigenvalues $e^{2\pi{\rm i}kq/N}$, $k\in\{0,1,\dots,N/q-1\}$. There $S^{-qr}$ is just an overall phase factor $e^{-2\pi{\rm i}qkr/N}$. Therefore $F_{q,r}$ is up-to this phase equal to an appropriate power of $\tilde{F}_{q,r}$.